OpenStudy (ray_sihota):

always have issues with fractions have problems with this equation. solve by substitution. 3x-2y=3 x-1/3y=8

OpenStudy (anonymous):

Solve one of the equations for Y.

OpenStudy (ray_sihota):

-2y=-3+2 divide by 2 y= 3/2+1

OpenStudy (anonymous):

Are the two equations you put in the original question the two in the system of equations?

OpenStudy (ray_sihota):

yes

OpenStudy (anonymous):

Because if so, the easiest way to do it is actually to solve for x in the second equation, then plug in what x is equal to into the first equation. You then solve for y like normal, at which point you should have the final value for y. Plug the number for y into either equation, and solve for x. You now have the values for x and y.

OpenStudy (anonymous):

If you need to use substitution, the easiest way to do it is to find an x or y value that has no coefficient, so things don't get complicated.

OpenStudy (anonymous):

Or technically, the coefficient is 1, since every number has some coefficient.

OpenStudy (ray_sihota):

if i have x-1/3y=8 = -1/3y =8-x

OpenStudy (ray_sihota):

and then substitute that into the first eqation

OpenStudy (ray_sihota):

how is the coefficient technically 1?

OpenStudy (kropot72):

\[3x-2y=3\ ............(1)\] \[x-\frac{1}{3}y=8\ ............(2)\] We can find a value for x in terms of y from equation (2) by adding (1/3)y to both sides, giving: \[x=\frac{1}{3}y+8\ ........(3)\] Now we can substitute the value of x in equation (3) into equation (1) giving: \[3(\frac{1}{3}y+8)-2y=3\ .......(4)\] Removing the brackets in equation (4) gives: \[y+24-2y=3\ ......(5)\] Now you should be able to solve equation (5) to find the value of y.

OpenStudy (ray_sihota):

thank you

OpenStudy (kropot72):

You're welcome :)